OpenStudy (anonymous):

-4x+2y=-8 how to put in slope intercept help asap

OpenStudy (texaschic101):

slope intercept form is y = mx + b -4x + 2y = -8 --- add 4x to both sides -4x + 4x + 2y = 4x - 8 -- combine like terms 2y = 4x - 8 -- divide both sides by 2 2y/2 = 4/2x - 8/2 y = 2x - 4 --- slope intercept form Basically, your just solving for y any questions ?